OpenStudy (lgbasallote):

for calculus enthusiasts... \(\large \int (\sec x + \csc x)^2 dx\) how to solve?

OpenStudy (anonymous):

Open it/ Expand it

OpenStudy (lgbasallote):

i got \(\large \cot x - 2\int \frac{1}{\sin x \cos x} + \tan x\)

OpenStudy (anonymous):

\[4\int \frac{1}{\sin 2x}\]

OpenStudy (anonymous):

\[4 \int \csc 2x \]

OpenStudy (lgbasallote):

\(\Large \mathtt{\color{pink}{uh} \color{lightgreen}{-} \color{lightblue}{huh}}\)

OpenStudy (anonymous):

\[4 \int\limits \csc^2 2x\quad dx+4\int\limits \csc2x \quad dx\]

OpenStudy (lgbasallote):

^where did that come from?

OpenStudy (amistre64):

i think cinar posted it, but i cant be too sure :P

OpenStudy (lgbasallote):

i cant see the post @_@

myininaya (myininaya):

\[(\sec(x)+\csc(x))^2=\sec^2(x)+2\sec(x)\csc(x)+\csc^2(x)\] \[=\sec^2(x)+2\frac{1}{\cos(x)}\frac{1}{\sin(x)}+\csc^2(x)\] \[=\sec^2(x)+\frac{2}{\cos(x) \sin(x)}+\csc^2(x)\] \[=\sec^2(x)+\frac{2 \cdot 2}{2 \cos(x) \sin(x)} +\csc^2(x)\] \[=\sec^2(x)+\frac{4}{\sin(2x)}+\csc^2(x)\] \[=\sec^2(x)+4 \csc(2x)+\csc^2(x)\]

OpenStudy (australopithecus):

this looks fun im too late though :(

OpenStudy (amistre64):

at 1245am, youre all too late

OpenStudy (amistre64):

or is that 0045 hours?

OpenStudy (australopithecus):

its whatever you want it to be

OpenStudy (amistre64):

0075 if we use military metric lol

OpenStudy (lgbasallote):

yes @myininaya the first and last terms are very much integrable...but the middle term i dont know how :C

OpenStudy (anonymous):

might have an error >.<

myininaya (myininaya):

\[\int\limits\limits_{}^{}\csc(2x) dx=\int\limits\limits_{}^{}\csc(2x) \cdot \frac{\csc(2x)+\cos(2x)}{\csc(2x)+\cot(2x)} dx \] Let \[u=\csc(2x)+\cot(2x) => du=-2 \csc(2x)(\cot(2x)+\csc(2x)) dx\]

OpenStudy (anonymous):

\[-2cosec2xcos2x-2\ln|cosec2x+\cot2x|+C\]

OpenStudy (lgbasallote):

hmm @cinar there's still that \(4\int \csc (2x)\) problem in your solution :p

myininaya (myininaya):

\[\int\limits_{}^{}\frac{du}{-2 u}=\frac{-1}{2} \ln|u|+C\]

OpenStudy (lgbasallote):

i dont get the multiplication thing you did myin :P im sure that wasnt conjugate...

OpenStudy (anonymous):

can't you use the table..

myininaya (myininaya):

for integrating sec(x) and csc(x) you need to multiply by a weird one

OpenStudy (anonymous):

my post was only for the:\[2\int\limits\sec x \csc xdx \]term after you square.

myininaya (myininaya):

for sec(x) it is (sec(x)+tan(x))/(sec(x)+tan(x)) and for csc(x) it is (csc(x)+cot(x))/(csc(x)+cot(x))

OpenStudy (anonymous):

That's quite neat @joemath314159. Always like tricks like that :-)

OpenStudy (anonymous):

ah, the signs are backwards though, ln(cos) should be the negative one.

OpenStudy (lgbasallote):

hmm seems @joemath314159 gave the most convincing and comprehensible reply lol..thanks :DDD

OpenStudy (lgbasallote):

though i think the one multiplied to csc x is cscx - cotx @myininaya i mean that's what i remember..is it +?